OpenStudy (anonymous):

A biologist has determined that a particular osprey has a 70% chance of catching a fish on any given day. Carry out a simulation of twenty trials using the random number table below to find the probability that the osprey will actually catch a fish on all of the next three days.

OpenStudy (anonymous):

945 025 354 793 236 106 746 981 105 012 832 180 250 871 835 793 726 864 496 947

OpenStudy (anonymous):

@radar

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@phi , @.Sam.

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

SOMEONE PLEASE HELP ME!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

@radar

OpenStudy (amistre64):

ive never used a random number table, ive seen them but havent really delved into how to use them. can you explain to me the process? since it should be covered in your material.

OpenStudy (amistre64):

i think the table will generate the yes-no trials; then 70%, times the number of (yes, yes, yes) outcomes in 20 would seem to be it

OpenStudy (anonymous):

Well my teacher said the answer was 35% but i just don't know how to do the work for it

OpenStudy (anonymous):

@radar @amistre64

OpenStudy (amistre64):

the key is in knowing how to work the table. spose the table gives us x/20 results that are y,y,y then .7*x/20 = .35 if the teacher is correct so x=10 ... the table would most likely produce about 10 instances

OpenStudy (amistre64):

since i have no real idea how to use the table to produce random numbers then the process is a mystery to me as how we find x :)

OpenStudy (anonymous):

@amistre64 could you help me with other probability questions?

OpenStudy (amistre64):

maybe, im not the greatest at them but for like stats 1 i can prolly get by

OpenStudy (anonymous):

This is Algebra 2, my teacher is teaching us probability

OpenStudy (amistre64):

oh, then most likely i can fake it :)

OpenStudy (anonymous):

Okay thank you!! Mana has a candle box containing 6 red candles and 6 white candles. What is the probability the she will pull out three red candles and two white candles?

OpenStudy (anonymous):

THis is what i have \[6P3(6P2)\div12P5=5/132\]

OpenStudy (amistre64):

the hardest part about this might be to consider all the ways that she can pull 5 candles such that 2 are white and 3 are red. algebra 2 i spose hasnt covered binomials

OpenStudy (amistre64):

oh good, youve got a little

OpenStudy (anonymous):

or \[6C3(6C2)\div12C5=225/792\]

OpenStudy (anonymous):

But idk which one

OpenStudy (amistre64):

if i brute it some to check: the ways to pull are: rrrww rrwrw rwrrw wrrrw rrwwr rwrwr wrrwr rwwrr wrwrr wwrrr so 10 different way to pull this combination, all we have to do is assign values to the pulls now

OpenStudy (amistre64):

rrrww r: 6 out of 12 r: 5 out of 11 r: 4 out of 10 w: 6 out of 9 w: 5 out of 8 agreed so far?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

lets pull another set: the out of parts will remain the same regardless so lets focus on the other part wrrwr: w:6 r: 6 r: 5 w:5 r: 4 agreed?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

if we do this for each one, we would see that this 66554 is just getting mixed around but the product of them never changes. 6*6*5*5*4 is the same regardless of the order, right?

OpenStudy (anonymous):

Right

OpenStudy (amistre64):

so:\[10*\frac{^6P_3~ ^6P_2}{^{12}P_5}\] is the way im seeing it

OpenStudy (amistre64):

or it may be 10 times out of 5! (120) total outcomes

OpenStudy (amistre64):

nah, its just the addition of all the possible success

OpenStudy (amistre64):

10.6.5.4.6.5 ----------- 12.11.10.9.8 .6.5.4.6.5 ----------- 12.11..9.8 ..5.5 ------ 6.11... 25/66 is what im getting 25/60

OpenStudy (amistre64):

lol, 25/60 was a priorcalc i forgot to delete

OpenStudy (anonymous):

I see what you did, but could you try the combination, permutation method to see?

OpenStudy (anonymous):

You still there @amistre64

OpenStudy (amistre64):

im here, sites getting laggy tho

OpenStudy (anonymous):

Okay, could you try the combination, permutation method to see?

OpenStudy (anonymous):

Because idk when to use combination or when to use permutation

OpenStudy (amistre64):

i have a hard time with it myself .... my teacher had us probing poker hands, it was horrible

OpenStudy (amistre64):

6C3 * 6C2 ---------- = 25/66 also 12C5 nCk = nPk/k! 6P3 * 6P2 * 5! ------------- 3! 2! 12P5 5.4.3.2.1 20 -------- = -- = 10 3.2.1.2.1 2

OpenStudy (amistre64):

soo ..\[\frac{6C3~6C2}{12C5}=10\cdot \frac{6P3~6P2}{12P5}\] :)

OpenStudy (amistre64):

the idea is simply to use what suits you best. I can reason better from the 10 cases of 'pick's, rather than the cases of 'choose'

OpenStudy (anonymous):

Did you do 6C3.... times 6P3.....?

OpenStudy (amistre64):

no, i just went over how the cases relate to one another. in essense, use P of you can count the number of times that set can occur ... and multiply the results by that number use C if you feel more comfortable overall

OpenStudy (anonymous):

I am so confused haha

OpenStudy (amistre64):

spose we wanted 4 red and 1 white rrrr w is the basic setup, there are 5C4 ways to arrange this, right?

OpenStudy (anonymous):

wouldnt it be 6C4? Since there are 6 red candles

OpenStudy (amistre64):

not the direction im headed, im simply trying to count the ways that we can have 5 objects arranged, such that 4 are the same and 1 is the same. there is a rule for this: n objects, such that a are the same, b are the same, c are the same, .... k are the same the number of ways that we can arrange them is:\[\frac{n!}{a!~b!~c!~...~k!}\] in this case we have 5 objects, 4 are the same and 1 is the same:\[\frac{5!}{4!~1!}=\frac{5!}{4!}=5C4\]

OpenStudy (anonymous):

Oh okay

OpenStudy (amistre64):

im pretty sure we can reason that the setups:\[\frac{6C3~6C2}{12C5}=\frac{5!}{3!~2!}\frac{6P3~6P2}{12P5}\] so if you want to use one or the other it would be fine

OpenStudy (amistre64):

if you use the P setup, it has a multiplier to deal with

OpenStudy (amistre64):

spose we have a different setup: 7 red, and 5 white; such that we want 3 red and 2 white ...

OpenStudy (anonymous):

okay...

OpenStudy (amistre64):

\[\frac{7C3~5C2}{12C5}\] \[\frac{5!~7P3~5P2}{3!~2!~12P5}\] \[\frac{7C3~5C2}{12C5}=\frac{5!}{3!~2!}\frac{7P3~5P2}{12P5}\]

OpenStudy (amistre64):

if we do the P setup, we still need to determine how many times it occurs: which is still 5C3 (or 5C2) = 10 times the C setup is just a 'simplification' of the P setup

OpenStudy (anonymous):

Why are you including 5! and 3! in the equation?

OpenStudy (amistre64):

spose we have 5 cats, 3 dogs, and 2 rabbits; 10 objects total what the probability of getting 3 cats, 1 dog, and 2 rabbits? 6 objects total \[\frac{6!}{3!~1!~2!}~\frac{5P3~3P1~2P2}{10P6}\]

OpenStudy (amistre64):

its a counting rule: mentioned above with n objects such that such and such are the same

OpenStudy (anonymous):

I guess i understand, but for my question above was my answer, 5/132 or 25/88

OpenStudy (amistre64):

you might want to recalculate those options you posted

OpenStudy (amistre64):

and, 5/132 is calced correctly 10*5/132 = 25/66 which is what i would expect of the Ps

OpenStudy (anonymous):

Oh okay thanks! So is it 25/66(combination) or 5/132(permutation)

OpenStudy (amistre64):

without the multiplier 10, yeah thats what you would get. to correct the permutations we would have actually use that process correctly :)